# Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges

### v1

@@
Abstract.
Creating engaging narratives from visual data is crucial for automated digital media consumption, assistive technologies, and interactive entertainment. This survey covers methodologies used in the generation of these narratives, focusing on their principles, strengths, and limitations.
The survey also covers tasks related to automatic story generation, such as image and video captioning, and visual question answering, as well as story generation without visual inputs. These tasks share common challenges with visual story generation and have served as inspiration for the techniques used in the field. We analyze the main datasets and evaluation metrics, providing a critical perspective on their limitations.
Visual Story Generation, Image Captioning, Visual Question Answering, Storytelling
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 00journalnumber: 0article: 000publicationmonth: 0ccs: Computing methodologies Natural language generationccs: Computing methodologies Video summarizationccs: Computing methodologies Scene understanding
1. Introduction
Stories are fundamental to human experience, serving as a bridge between imagination and reality (Hühn et al., 2014). They represent the art of telling tales, whether real or imagined, and present a sequence of events that can captivate, inform, and provoke thought. Storytelling has been an essential part of human culture for millennia, serving as a means of communication, entertainment, education, and moral guidance. The art of storytelling has evolved over time, beginning with the practice of narrating stories through spoken words, advancing to written literature, and, more recently, transitioning to digital media formats. Stories have played a crucial role in shaping societies, passing down knowledge and values from one generation to another. Whether told around a campfire, written in books, or shared through videos and interactive platforms, storytelling continues to captivate and connect people, transcending time and technology (Abbott, 2008).
The field of automated story generation has gained attention, fueled by advances in Natural Language Processing (NLP) (Jurafsky and Martin, 2009) and computer vision (Szeliski, 2010). The goal is to create systems capable of generating novel and engaging stories, similar to those produced by humans (Gervas, 2009). With the increasing availability of images and videos, there is a growing interest in generating stories based on this visual data (Krishna et al., 2017). We refer to this task as Visual Story Generation (VSG).
1.1. Definition and challenges
@@
  • Understanding visual content: The recognition of objects and scenes in the visual data, as well as understanding the relationships between them is crucial for generating coherent and meaningful narratives (Xu et al., 2015).
@@
1.2. Background and Motivation
The motivations for VSG ] are numerous. For example, VSG can provide an interactive component to media consumption. Users could potentially influence the direction of the story by choosing different sequences of images or videos, creating a unique, personalized narrative. This could be particularly beneficial in areas like interactive gaming and movies, educational platforms, or even in digital marketing, where personalized narratives can greatly enhance user engagement and satisfaction. VSG can also be used in journalism: it can be used to summarize large amounts of visual data, giving insights into the underlying patterns and relationships, and assisting in content creation.
The growing interest in this area is also motivated by its potential to contribute to the understanding of human cognition and creativity. By developing models that emulate human-like storytelling abilities, researchers can gain insights into the cognitive processes involved in story creation, as well as explore the underlying mechanisms of creativity. These approaches can further lead to novel applications, such as more engaging virtual experiences (Fan et al., 2018; Peng et al., 2018).
@@ 1.2. Background and Motivation
1.3. Scope and structure of the survey
We start by defining the core elements of a story: they are essential for understanding the structure and dynamics of a story and served as foundational components for the development of the early story generation systems. Section 2 presents these elements and discusses their significance in storytelling. It also clarifies the distinctions between story, narrative, and plot, illustrating how a single story can be narrated in multiple ways to yield diverse experiences.
We then explore the main datasets used in story generation, and the various metrics used to assess their performance. We address the limitations that current evaluation metrics face in such an open-ended task and discuss potential directions for future research in this regard. Section 3 addresses these issues.
Since they form the background for VSG approaches, we also explore the methods for textual story generation: from the early rule-based systems to deep learning models. We focus on deep learning methodologies, especially transformer architectures and Large Language Models (LLMs) (Bolaños et al., 2015; Ammanabrolu et al., 2019), due to their state-of-the-art performance. Section 4 presents these approaches and emphasizes their contributions and historical significance.
@@ 1.3. Scope and structure of the survey
Finally, Section 7 presents the key findings of this survey and proposes directions for future research.
2. The Elements of Stories
The essence of a story lies in its ability to connect a series of events and experiences, often conveyed by a narrator to an audience. These can range from personal intimate stories to larger epics that span cultures and time.
A story can be told in multiple ways using different narratives. These narratives are more than mere descriptions or dramatic enactments. They involve a process of narration where events are selected, arranged, and em-plotted to give meaning to experiences. This structuring of events is what forms the plot or the story’s synopsis.
The terms story, narrative, and plot are often used interchangeably in everyday language, but they have distinct meanings. According to the Oxford English Dictionary (Simpson, 2002) a story is the straightforward, chronological sequence of events, encompassing characters, settings, and actions. The story is the “what” – what happened, who it happened to, where and when it happened. In contrast, the narrative shapes the form or the “how”, adding context, interpretation, perspective, tone, and style of the storytelling, and nuance to the same set of events. The plot is connected to the narrative, referring to the sequence of events as they are presented to the audience. Thus, a single story can be narrated in multiple ways, each having a different plot and offering a distinct experience to the audience.
Narratives can be broadly categorized as linear or nonlinear (Lau and Chen, 2008), depending on their structure. Each type offers a different approach to unfolding a story and has distinct implications for how the audience experiences and interprets it.
Linear narratives follow a straightforward chronological progression, where events unfold in a sequential and cause-effect manner. Linear narratives often have a single protagonist or a group of characters whose actions and decisions drive the plot forward. This structure is commonly used in traditional storytelling formats. Linear narratives provide a sense of progression and closure, making them easy to follow and understand.
Nonlinear narratives deviate from the traditional linear structure by presenting events out of chronological order. They often involve flashbacks, flash-forwards, or parallel storylines that disrupt the temporal sequence. These narratives can create intrigue, suspense, and complexity by challenging the audience’s expectations and requiring them to piece together the fragmented storyline. The film “Cloud Atlas” (Mitchell, 2008) is an example.
@@ 2. The Elements of Stories
Story elements – character, conflict, theme, setting, plot, and mode – are the essential components that form the foundation of a story and are vital for both crafting and analyzing stories. In the following subsections, we detail each of these elements, discussing their significance in shaping the story.
2.1. Character
Characters are the entities through which the story unfolds (DiBattista, 2011). Their decisions and actions are pivotal in propelling the plot forward. Characters are typically people or other beings, fictional or based on a real-life person, and through their actions and speech, they often convey essential motives and emotions.
The way an audience initially perceives a character significantly influences their ongoing perception and emotional connection, such as empathy or antipathy. This initial impression sets the stage for the audience’s expectations regarding the future behavior of the character. Characters who deviate unexpectedly from their established patterns of behavior can create confusion or surprise, challenging the audience’s understanding and expectations.
Narratives usually center around main characters or protagonists. They are the focal points of the story, often encountering central conflicts, undergoing significant personal growth, or acquiring new knowledge. Many stories also have antagonists, adding complexity and tension to the storyline. In traditional stories, the protagonist often takes the form of a hero, a sympathetic figure battling for morally good causes. Conversely, the antagonist may be portrayed as a villain, acting against these good causes or even perpetrating evil.
Narratives often employ a diverse range of character types, each contributing uniquely to the story’s depth and dynamics. The way characters are developed, presented, and evolve throughout the narrative plays a crucial role in engaging the audience and driving the story’s progression.
2.2. Conflict
Conflict is a fundamental element that generates tension and drives the thoughts and actions of characters (Abbott, 2008). The manner in which conflicts are introduced, developed, and resolved plays a significant role in shaping the trajectory of the narrative and the engagement of the audience with the story. A well-crafted conflict not only propels the narrative forward but also enriches the emotional and psychological depth of the characters and the story itself. The resolution of the conflict, typically towards the end of the narrative, is a crucial aspect. It marks the culmination of tension and provides closure to the story’s central problem.
Broadly, conflict encompasses any form of tension that influences characters, while more narrowly, it may refer to the main challenge faced by the protagonist of the story. This primary conflict often leads to additional struggles for the protagonist, such as feelings of anxiety, insecurity, indecisiveness, or other mental difficulties. These struggles can be categorized as secondary or internal conflicts, adding layers to the journey of the character and depth to the story. Longer stories typically involve multiple conflicts. These can be smaller-scale issues that occur alongside or as part of the main conflict, contributing to the complexity of the story. Commonly used conflict types include: “Character vs. Character” (e.g., opposition); “Character vs. Nature” (e.g., natural forces); “Character vs. Society” (e.g., societal norms); “Character vs. Unavoidable Circumstances” (e.g., fate); and “Character vs. Self” (e.g., dilemmas).
2.3. Theme
Themes encompass the major underlying ideas presented by a story, often abstract and open to the audience’s interpretation (Griffith, 2010). Unlike more concrete story elements, themes invite discussion and contemplation, allowing for a range of interpretations and conclusions. They are not definitive statements but rather exploratory concepts that encourage the audience to engage with and reflect on the story’s deeper meanings. As the narrative unfolds, themes can evolve or reveal new dimensions, offering the audience opportunity to develop their own ideas about the messages or insights the story conveys. The subjective nature of themes means that different audience members may arrive at varying interpretations, which might differ from or even go beyond the creator’s original intent. Thus, understanding and incorporating themes in narrative generation allows for a dynamic, thought-provoking storytelling experience that resonates with and challenges the audience. An example of a theme is the elusive nature of the American Dream, as portrayed by F. Scott Fitzgerald in “The Great Gatsby” (Fitzgerald and Reynolds, 1993).
2.4. Setting
@@
The setting is an element that encompasses the time, place, and context in which the story unfolds (Truby, 2007). It includes the physical and temporal surroundings inhabited by the characters, as well as the social or cultural conventions that influence their actions and decisions. The setting can act dynamically, almost like a character itself, possessing specific traits and undergoing changes that directly affect the plot and contribute to the story’s development. This dynamic nature of the setting allows it to evolve over the course of the story, reflecting and impacting the narrative’s progression. The Hogwarts School of Witchcraft and Wizardry in the “Harry Potter” series (Rowling, 2015) is an example of an evolving setting that influences both plot and character development and impacts the progression of the story.
2.5. Plot
@@
The plot is the sequence of events, the plot points, that propels the story from beginning to middle to end (Dibell, 1999). Each plot point is a moment of change that influences the understanding, decisions, and actions of the characters. This progression is often closely tied to the protagonists encountering or realizing a conflict, and their subsequent efforts to resolve it. This journey creates emotional stakes for both the characters and the audience, generating a sense of suspense. The plot’s movement is integral to the narrative, as it not only advances the storyline but also deepens the audience’s emotional investment in the fate of the characters. A single story can have multiple plots, each representing a different sequence of events or perspective. The plot is usually structured through cause and effect, where actions of the characters and events lead to reactions and consequences, forming a chain of scenes that collectively move the story forward.
As an example, in “The Lord of the Rings” by J.R.R. Tolkien (Tolkien, 2012), the plot unfolds as a series of events beginning with Frodo inheriting the One Ring and learning about its power and dangers, followed by, among other events, the Council of Elrond and the formation of the Fellowship, Gandalf’s fall, the breaking of the Fellowship, and Frodo’s decision to, at first, continue alone. These plot points are influenced by character decisions and actions, which lead to various developments and challenges, enhancing the narrative.
2.6. Mode
The mode is a set of choices and techniques employed by the author to frame the narrative (Pinault, 1992). A single story can be narrated using different modes, each offering a different perspective or experience to the audience. The mode includes the scope of the presented information, the style and type of language used, the medium through which the story is conveyed, and the manner and extent of narrative exposition and commentary.
@@ 2.6. Mode
A key component of the mode is the point of view or perspective from which the story is told. A first-person narrative is an example where a character – often the protagonist – narrates the story using personal pronouns like “I” and “me” directly engaging the audience with their perspective. A third-person narrative is another example that avoids personal pronouns, offering a more detached or impersonal viewpoint. These perspectives are examples of narration, a technique that involves the explicit telling of a story through spoken or written commentary. While narration is frequent in written narratives, it can be optional in other forms, like visual or interactive media.
3. Benchmark Datasets and Evaluation Metrics
This section presents and discusses benchmark datasets used for story generation and evaluation metrics that assess the performance of story generation models.
Although not all approaches are data-driven, some of them, especially deep learning models and other machine learning approaches, rely heavily on data to guide the learning process. Various story datasets have been employed in the development of story generation systems by the community.
@@ 3. Benchmark Datasets and Evaluation Metrics
Table 1 provides a compilation of the datasets documented in the literature for both VSG and textual story generation. Among these datasets, VIST (Huang et al., 2016) is prominently utilized by the research community. This dataset serves as a primary benchmark, and most publications in this field report their experimental results on it. The Visual Writing Prompts dataset (Hong et al., 2023) is noteworthy for its focus on stories that are visually grounded on the characters and objects present in the images. Similarly, the Writing Prompts dataset (Fan et al., 2018) and STORIUM (Akoury et al., 2020) are important because they contain long stories derived from the imagination of the users of the platforms where they were collected.
These datasets allow for the exploration of more complex and imaginative narrative structures, serving as a resource to train models to generate fictional stories.
Dataset No. of Stories Type Notes
NY City (Park and Kim, 2015) 11,863 visual 11,863 blog posts / 78,467 images
Disneyland (Park and Kim, 2015) 7,717 visual 5 7,717 blog posts / 60,545 images
ROCStories (Mostafazadeh et al., 2016) 100,000 visual 5 sentences/story.
LSMDC (Rohrbach et al., 2016) 118,114 visual 202 movies.
VIST (Huang et al., 2016) 50,000 visual 210,819 unique photos
@@
Visual Writing Prompts (Hong et al., 2023) 12,000 visual 5 to 10 images/story, grounded on the characters.
Pororo-SV (Kim et al., 2017) 27,328 visual
YouCook2 (Das et al., 2013) 2,000 visual Average length of 5.26 mins.
Writing Prompts (Fan et al., 2018) 300,000 text 28.4 words/prompt, 734.5 words/story.
@@
CMU movie summary corpus 42,306 visual Includes box office revenue, genre, release date, runtime, character names and actor names.
Wikipedia’s movie plots 34,886 text Summary of the movie plot with metadata.
Sci-fi TV plot summaries (Ammanabrolu et al., 2019) 2,276 text 80+ episodes of TV Series.
STORIUM (Akoury et al., 2020) 5,743 text 19,278 words/story. Annotations for scene entries and cards played to guide the story.
Table 1. Story Generation Datasets.
The evaluation metrics, used to assess the performance of VSG models, provide quantitative measures to gauge the quality and coherence of generated narratives. Most of these metrics were originally developed for machine translation or summarization but have been adopted by the community for other tasks such as image captioning and story generation: instead of comparing a generated text with a reference translation, the generated text is compared with a reference caption in the case of image captioning, or with a reference story in the case of story generation. In the following subsections, we discuss commonly used metrics, including their strengths and limitations.
3.1. Perplexity
Perplexity (Jurafsky and Martin, 2023) evaluates the performance of autoregressive language models. It measures the model’s ability to predict a sequence and is defined as the exponential average negative log-likelihood of a sequence: given a tokenized sequence $X=(x_{1},x_{2},\ldots,x_{N})$, where $N$ is the length of the sequence, the perplexity of $X$ is given by Eq. 1. $P(x_{i}|x_{1},\ldots,x_{i-1})$ denotes the probability of token $x_{i}$ given the preceding token sequence $x_{1},\ldots,x_{i-1}$ as predicted by the model. Perplexity can be interpreted as the geometric mean of the inverse probability of the word sequence, indicating how well the model predicts each subsequent token in the sequence. Lower values indicate better predictive performance, meaning that the model is more effective in capturing the underlying structure of the language.
(1) $\text{Perplexity}(X)=\sqrt[N]{\prod_{i=1}^{N}\frac{1}{P(x_{i}|x_{1},\ldots,x_{i-1})}},$
3.2. BLEU
Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) evaluates the quality of generated text by comparing it with one or more reference texts. It was developed for machine translation tasks, but has been adopted for various natural language generation tasks including VSG. BLEU outputs a score between 00 and $1$, indicating the similarity between the generated and reference texts. Higher scores reflect better alignment with the reference: a score of $1$ signifies perfect agreement.
BLEU compares n-grams in the candidate text with the n-grams in the reference texts, computing the precision of these matches, which are independent of their position in the text. The general idea is that the greater the number of matching n-grams, the higher the quality of the text. To mitigate the issue of favoring shorter texts (since they can potentially have higher precision), BLEU incorporates a brevity penalty. This penalty balances precision and recall, ensuring that the metric does not disproportionately favor shorter stories.
In Eq. 2, BP is the brevity penalty (Eq. 3), $w_{n}$ is the weight for each n-gram, $p_{n}$ is the n-gram precision, and $N$ is typically up to four, accounting for the use of unigrams, bigrams, trigrams, and four-grams.
(2) BLEU $=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$
(3) BP =$=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases}$
While widely used, BLEU has limitations in evaluating stories involving word variations, synonyms, and creative text. Its reliance on exact n-gram matches fails to capture semantic equivalence when different wording or phrasing is used, potentially penalizing stories that are semantically accurate but stylistically different. This results in reduced effectiveness in assessing linguistic richness, idiomatic expressions, and the subtleties of creative language use.
3.3. Self-BLEU
Self-BLEU (Zhu et al., 2018) evaluates diversity within a set of generated data. While the original BLEU was designed to assess the similarity between two texts, Self-BLEU adapts this concept to measure how much each sentence within a generated text resembles the rest of the text. An individual sentence is treated as the hypothesis and all other companion sentences are considered as the reference set. The BLEU score is calculated for each sentence against the aggregated reference set, and the average of these scores constitutes the Self-BLEU score for the entire document. In Eq. 4, $N$ is the number of sentences, $s_{i}$ is the $i$-th sentence, and $S\setminus\{s_{i}\}$ denotes the set of all sentences excluding $s_{i}$.
[…]
A higher Self-BLEU score indicates a lower diversity among the generated sentences, suggesting a potential mode collapse issue within the generative model. Mode collapse occurs when the model produces a narrow range of outputs, repeatedly generating similar types of data while failing to capture the full diversity of the target distribution. Conversely, a lower Self-BLEU score suggests greater diversity and a healthier, more varied generation process. This is particularly relevant for story generation models where diversity in narrative output is crucial for creating engaging and unique stories. High Self-BLEU scores might signal a lack of creativity and variability in the stories produced, whereas lower scores indicate a model’s effectiveness in generating rich, varied narratives.
3.4. METEOR
Metric for Evaluation of Translation with Explicit ORdering (METEOR) (Banerjee and Lavie, 2005) was designed to evaluate machine translation outputs, addressing some limitations of BLEU. It focuses on sentence-level evaluation, aiming to achieve higher correlation with human judgments. The metric relies on the harmonic mean of unigram precision and recall, with emphasis placed on recall to better align with human evaluative tendencies.
The algorithm initiates by aligning the candidate text with the reference text, focusing on unigram mappings. These mappings are constrained such that each unigram in the candidate text aligns with zero or one unigram in the reference text. The alignment process prioritizes configurations with the fewest crossing lines to ensure clarity and coherence.
Precision $P={m}/{w_{t}}$ and recall $R={m}/{w_{r}}$ are the ratios between $m$ (the number of matching unigrams in the candidate text) and either $w_{t}$ (the number of unigrams in the candidate text) or $w_{r}$ (the number of unigrams in the reference text). These metrics are integrated using the harmonic mean, emphasizing recall $F_{mean}=10\cdot P\cdot R/(R+9\cdot P)$. To account for the arrangement of words, METEOR introduces a penalty $p=0.5(c/u_{m})^{3}$, where $c$ is the number of chunks, and $u_{m}$ is the number of mapped unigrams. The final METEOR score for a candidate text is given by $M=(1-p)F_{mean}$.
3.5. CIDEr
Consensus-based Image Description Evaluation (CIDEr) (Vedantam et al., 2014) was developed for image captioning evaluation. It evaluates how well a candidate sentence $c_{i}$ aligns with a consensus among a set of reference sentences $S_{i}=\{s_{i1},\ldots,s_{im}\}$ for an image $I_{i}$. CIDEr involves stemming words in the sentences to their root forms. Each sentence is represented by its n-grams, where an n-gram $\omega_{k}$ is an ordered set of words, and n-grams from one to four words are considered.
CIDEr evaluates the consensus by measuring how frequently n-grams in the candidate sentence appear in the reference sentences. It also assigns lower weights to n-grams common across all images in the dataset, as they are likely less informative. The Term Frequency-Inverse Document Frequency (TF-IDF) (Leskovec et al., 2014) method is employed to weight each n-gram $\omega_{k}$ based on its informativeness, as shown in Eq. 5, where $h_{k}(s_{ij})$ is the frequency of $\omega_{k}$ in sentence $s_{ij}$, and $\Omega$ is the set of all n-grams and $|I|$ is the set of all images in the dataset.
(5) $g_{k}(s_{ij})=\frac{h_{k}(s_{ij})}{\sum_{\omega_{l}\in\Omega}h_{l}(s_{ij})}\log\left(\frac{|I|}{\sum_{I_{p}\in I}\min(1,\sum_{q}h_{k}(s_{pq}))}\right),$
The CIDEr score for n-grams of length $n$ is calculated using the average cosine similarity between the TF-IDF vectors of the candidate and reference sentences, capturing both precision and recall as shown in Eq. 6, where $g_{n}(c_{i})$ and $g_{n}(s_{ij})$ are the TF-IDF vectors for n-grams of length $n$ in the candidate and reference sentences, respectively. $m$ is the number of reference sentences for the candidate image $i$.
(6) $CIDEr_{n}(c_{i},S_{i})=\frac{1}{m}\sum_{j}\frac{g_{n}(c_{i})\cdot g_{n}(s_{ij})}{\|g_{n}(c_{i})\|\|g_{n}(s_{ij})\|},$
To derive the final score, CIDEr combines scores across different n-gram lengths, as shown in Eq. 7, where uniform weights $w_{n}=1/N$ have empirically been found to perform best, and $N$ is typically set to 4.
(7) $CIDEr(c_{i},S_{i})=\sum_{n=1}^{N}w_{n}CIDEr_{n}(c_{i},S_{i}),$
3.6. ROUGE
Recall-Oriented Understudy for Gisting Evaluation (ROUGE) (Lin, 2004) is a recall-oriented measure of the overlap of n-grams between a candidate text and reference texts. It was developed for summarization tasks, but has been used in other natural language generation tasks. ROUGE-N, one of the most common variations of the metric, is the ratio of the number of n-grams co-occurring in the candidate text and the reference texts to the total number of n-grams in the reference texts, as shown in Eq. 8, where $\text{gram}_{n}$ is the n-gram being considered, and $\text{Count}_{\text{match}}(\text{gram}_{n}(S))$ is the maximum number of n-grams co-occurring in a candidate and a set of reference texts.
[…]
As the number of reference texts increases, the denominator in the ROUGE-N formula also increases, which is intuitive since adding more references expands the space of acceptable texts, enhancing the robustness of the evaluation.
For multiple references, ROUGE-N is calculated as the maximum of the pairwise ROUGE-N scores between the candidate text and each reference text (Eq. 9, where $r_{i}$ is a reference text, and $s$ is the candidate text).
[…]
This approach is also extended to other ROUGE metrics such as ROUGE-L (based on the longest common subsequence), ROUGE-W (considers word sequences with gaps), and ROUGE-S (based on co-occurrence statistics of skip-bigrams, i.e., pairs of words in sentence order, allowing arbitrary gaps between them).
3.7. BERTScore
BERTScore (Zhang et al., 2019) is a metric that uses contextual embeddings to evaluate the similarity between a reference sentence $x=\{x_{1},\ldots,x_{k}\}$ and a candidate sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$. In story generation, $x$ and $\hat{x}$ are the reference and candidate stories, respectively. Each token in $x$ and $\hat{x}$ is a contextual embedding obtained from models trained on language modeling objectives, enabling the generation of distinct vectors for the same word in different contexts. The similarity between tokens is computed using cosine similarity. For pre-normalized vectors, this calculation simplifies to the inner product of the embeddings, that is, $\text{similarity}(x_{i},\hat{x}_{j})=x_{i}^{\top}\hat{x}_{j}$.
BERTScore computes recall $R_{\text{BERT}}$, precision $P_{\text{BERT}}$, and F1 score $F_{\text{BERT}}$ by matching each token in one sentence to the most similar token in the other sentence. These scores are defined as shown in Eqs. 1011, and 12, respectively.
(10) $R_{\text{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}x_{i}^{\top}\hat{x}_{j}$
(11) $P_{\text{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}x_{i}^{\top}\hat{x}_{j}$
(12) $F_{\text{BERT}}=\frac{2\cdot P_{\text{BERT}}\cdot R_{\text{BERT}}}{P_{\text{BERT}}+R_{\text{BERT}}}$
BERTScore can incorporate importance weighting using Inverse Document Frequency (IDF) scores, emphasizing rare words which are often more indicative of sentence similarity. Eq. 13 shows the recall with IDF weighting.
(13) $R_{\text{BERT}}^{\text{idf}}=\frac{\sum_{x_{i}\in x}\text{idf}(x_{i})\max_{\hat{x}_{j}\in\hat{x}}x_{i}^{\top}\hat{x}_{j}}{\sum_{x_{i}\in x}\text{idf}(x_{i})}$
To make BERTScore more interpretable, it is rescaled against an empirical lower bound $b$, derived from averaging BERTScores computed on randomly paired sentences from a large corpus. Eq. 14 shows the rescaled BERTScore $\hat{R}_{\text{BERT}}$.
(14) $\hat{R}_{\text{BERT}}=\frac{R_{\text{BERT}}-b}{1-b}$
3.8. Discussion

### v2

# Visual Story Generation; Image Captioning; Visual Question Answering; Storytelling

@@
Abstract
Creating engaging narratives from visual data is crucial for automated digital media consumption, assistive technologies, and interactive entertainment. This survey covers methodologies used in the generation of these narratives, focusing on their principles, strengths, and limitations. The survey also covers tasks related to automatic story generation, such as image and video captioning, and visual question answering. These tasks share common challenges with Visual Story Generation (VSG) and have served as inspiration for the techniques used in the field. We analyze the main datasets and evaluation metrics, providing a critical perspective on their limitations.
keywords:
Visual Story Generation; Image Captioning; Visual Question Answering; Storytelling
[…]
1 Introduction
In this document we survey the field of story generation from visual inputs, covering techniques, related tasks, and challenges. Stories are fundamental to human experience, serving as a bridge between imagination and reality Hühn et al. (2014). They represent the art of telling tales, whether real or imagined, and present a sequence of events that can captivate, inform, and provoke thought. Storytelling has been an essential part of human culture for millennia, serving as a means of communication, entertainment, education, and moral guidance. The art of storytelling has evolved over time, beginning with the practice of narrating stories through spoken words, advancing to written literature, and, more recently, transitioning to digital media formats. Stories have played a crucial role in shaping societies, passing down knowledge and values from one generation to another. Whether told around a campfire, written in books, or shared through videos and interactive platforms, storytelling continues to captivate and connect people, transcending time and technology Abbott (2008). The field of story generation has gained attention, fueled by advances in Natural Language Processing (NLP) Jurafsky and Martin (2009) and computer vision Szeliski (2010). The goal is to create systems capable of generating novel and engaging stories, similar to those produced by humans Gervas (2009). With the increasing availability of images and videos, there is a growing interest in generating stories based on this visual data Krishna et al. (2017). We refer to this task as VSG.
1.1 Definition and challenges
@@
  • Understanding visual content: The recognition of objects and scenes in the visual data, as well as understanding the relationships between them is crucial for generating coherent and meaningful narratives Xu et al. (2015).
@@
1.2 Background and Motivation
The motivations for VSG are numerous. For example, VSG can provide an interactive component to media consumption. Users could potentially influence the direction of the story by choosing different sequences of images or videos, creating a unique, personalized narrative. This could be particularly beneficial in areas like interactive gaming and movies, educational platforms, or even in digital marketing, where personalized narratives can greatly enhance user engagement and satisfaction. VSG can also be used in journalism: it can be used to summarize large amounts of visual data, giving insights into the underlying patterns and relationships, and assisting in content creation.
The growing interest in this area is also motivated by its potential to contribute to the understanding of human cognition and creativity. By developing models that emulate human-like storytelling abilities, researchers can gain insights into the cognitive processes involved in story creation, as well as explore the underlying mechanisms of creativity. These approaches can further lead to novel applications, such as more engaging virtual experiences Fan et al. (2018); Peng et al. (2018).
@@ 1.2 Background and Motivation
1.3 Scope and structure of the survey
We start by defining the core elements of a story in Section 2 as they are essential for understanding the structure and dynamics of a story. Section 2 presents these elements and discusses their significance in storytelling. It also clarifies the distinctions between story, narrative, and plot, illustrating how a single story can be narrated in multiple ways to yield diverse experiences.
In Section 3 we then explore the main datasets used in story generation, and the various metrics used to assess their performance. We address the limitations that current evaluation metrics face in such an open-ended task and discuss potential directions for future research in this regard.
@@ 1.3 Scope and structure of the survey
Section 6 examines the real-world applications of VSG. Finally, Section 7 presents the key findings of this survey and proposes directions for future research.
2 The Elements of Stories
Stories connect events and experiences through narration, ranging from personal tales to cultural epics. While the terms story, narrative, and plot are often used interchangeably, they have distinct meanings: a story is the chronological sequence of events (the “what”), the narrative is how these events are presented (the “how”), and the plot is the specific sequence as experienced by the audience Simpson (2002).
A single story can be told through multiple narratives, each offering different perspectives and experiences. Narratives may follow linear structures with chronological progression, or nonlinear approaches using flashbacks, flash-forwards, or parallel storylines that create complexity and require audience interpretation Lau and Chen (2008); Mitchell (2008).
@@ 2 The Elements of Stories
Story elements – character, conflict, theme, setting, plot, and mode – form the foundation for crafting and analyzing narratives. Understanding these elements is crucial for both human storytellers and automated story generation systems. In the following subsections, we detail each element and its significance in VSG.
2.1 Character
Characters are the entities through which stories unfold, with their decisions and actions propelling the plot forward DiBattista (2011). Through actions and speech, characters convey motives and emotions that shape audience perception and emotional connection. Initial character impressions influence audience expectations about future behavior. Characters who deviate from established patterns can create surprise or confusion, challenging audience understanding. Stories center around protagonists who encounter central conflicts, undergo personal growth, or acquire new knowledge. Antagonists add complexity and tension, often opposing the protagonist’s goals. Traditional narratives feature heroes battling for moral causes against villains who perpetrate evil. Character development and evolution throughout the narrative engages audiences and drives story progression. The range of character types contributes to story depth and dynamics.
2.2 Conflict
Conflict generates tension and drives character actions, shaping narrative trajectory and audience engagement Abbott (2008). How conflicts are introduced, developed, and resolved determines story progression and emotional depth. Conflict resolution provides closure to the central problem and marks the culmination of narrative tension. Conflict encompasses any tension influencing characters, from the protagonist’s main challenge to secondary internal struggles like anxiety or indecisiveness. These layers add complexity to character journeys and story depth. Longer narratives involve multiple conflicts that occur alongside the primary tension. Common conflict types include: Character vs. Character (opposition), Character vs. Nature (natural forces), Character vs. Society (social norms), Character vs. Unavoidable Circumstances (fate), and Character vs. Self (internal dilemmas).
2.3 Theme
Themes are the major underlying ideas of a story, often abstract and open to interpretation Griffith (2010). Unlike concrete story elements, themes invite discussion and allow for multiple interpretations and conclusions. They function as exploratory concepts that encourage audience reflection on deeper meanings.
Themes can evolve as narratives unfold, revealing new dimensions and allowing audiences to develop personal interpretations of the story’s messages. Different audience members may reach varying conclusions that differ from or exceed the creator’s original intent. Understanding themes in narrative generation creates dynamic, thought-provoking storytelling experiences.
2.4 Setting
@@
Setting encompasses the time, place, and context in which the story unfolds Truby (2007). It includes physical and temporal surroundings, as well as social or cultural conventions that influence character actions and decisions. Setting can act dynamically, possessing specific traits and undergoing changes that affect the plot. This allows it to evolve throughout the story, reflecting and impacting narrative progression. Hogwarts School in the “Harry Potter” series exemplifies an evolving setting that influences plot and character development Rowling (2015).
2.5 Plot
@@
Plot is the sequence of events that propels the story from beginning to end Dibell (1999). Each plot point represents a moment of change that influences character understanding, decisions, and actions. This progression typically involves protagonists encountering conflict and attempting resolution, creating emotional stakes and suspense for both characters and audience. Plot advancement deepens audience investment in character fate. Stories can have multiple plots representing different event sequences or perspectives. Plot structure follows cause and effect, where character actions and events lead to reactions and consequences, forming a chain that moves the story forward.
2.6 Mode
@@
Mode encompasses the choices and techniques authors use to frame narratives Pinault (1992). A single story can be narrated using different modes, each offering distinct audience perspectives and experiences. Mode includes information scope, language style, medium, and the extent of narrative exposition and commentary. A key component is point of view or perspective. First-person narratives use personal pronouns like “I” and “me,” directly engaging audiences with the character’s perspective. Third-person narratives avoid personal pronouns, offering more detached viewpoints. These perspectives exemplify narration techniques that explicitly tell stories through spoken or written commentary. While narration is common in written narratives, it can be optional in visual or interactive media.
3 Benchmark Datasets and Evaluation Metrics
@@
This section presents and discusses benchmark datasets used for story generation and evaluation metrics that assess the performance of story generation models. Table 1 provides a compilation of the datasets documented in the literature for VSG. Among these datasets, VIST Huang et al. (2016) is prominently utilized by the research community. This dataset serves as a primary benchmark, and most publications in this field report their experimental results on it. The VIST-Edit dataset Hsu et al. (2019) provides human edited versions of machine generated stories, offering insights into human preferences for story quality. The Visual Writing Prompts dataset Hong et al. (2023) is noteworthy for its focus on stories that are visually grounded on the characters and objects present in the images. These datasets allow for the exploration of more complex and imaginative narrative structures, serving as a resource to train models to generate fictional stories.
Dataset No. of Stories Notes
NY City Park and Kim (2015) 11,863 11,863 blog posts / 78,467 images
Disneyland Park and Kim (2015) 7,717 5 7,717 blog posts / 60,545 images
ROCStories Mostafazadeh et al. (2016) 100,000 5 sentences/story.
LSMDC Rohrbach et al. (2016) 118,114 202 movies.
VIST Huang et al. (2016) 50,000 210,819 unique photos
VIST-Edit Hsu et al. (2019) 14,905 Human-edited versions of 2,981 machine-generated stories
@@
Visual Writing Prompts Hong et al. (2023) 12,000 5 to 10 images/story, grounded on the characters.
Pororo-SV Kim et al. (2017) 27,328
YouCook2 Das et al. (2013) 2,000 Average length of 5.26 mins.
@@
CMU movie summary corpus 42,306 Includes box office revenue, genre, release date, runtime, character names and actor names.
Table 1: Visual Story Generation Datasets.